OpenStudy (kainui):

perl's sum question:

OpenStudy (kainui):

\[\LARGE \sum_{n=1}^{2000} (\frac{1}{ \sum_{k=1}^n k!})\]

OpenStudy (kainui):

Is there a closed form for this?

OpenStudy (perl):

right

OpenStudy (kainui):

I don't know, can you plug it into wolfram alpha like that?

OpenStudy (perl):

i did before , but now i forgot the syntax

OpenStudy (perl):

actually I used maple

OpenStudy (perl):

let me screen shot it

OpenStudy (perl):

I got 1.482622363

OpenStudy (kainui):

In simpler terms for other people who are curious, it looks like this: \[\Large \frac{1}{1!}+ \frac{1}{1!+2!}+ \frac{1}{1!+2!+3!}+...+ \frac{1}{1!+...+2000!}\]

OpenStudy (perl):

right

OpenStudy (kainui):

Slightly larger than e/2

OpenStudy (kainui):

And of course, smaller than e.

OpenStudy (perl):

first few sums produce the bulk of the sum, then it goes to zero fast

OpenStudy (kainui):

I wonder if in the limit as n goes to infinity if it approaches some sort of thing related to e?

OpenStudy (perl):

if someone could put that into wolfram, that would be nice :)

OpenStudy (perl):

is this considered a double series?

OpenStudy (perl):

a double series is Sum (Sum ( ... ) )

OpenStudy (kainui):

I don't think it is, at least not in this form it isn't.

OpenStudy (anonymous):

wow sum (k!) is divisible by 3 . just saying

OpenStudy (perl):

oh we might as well ask, does this converge as n goes to infinity?

OpenStudy (perl):

I think so

OpenStudy (kainui):

Definitely, since this is the limit for e: \[\LARGE e= \sum_{n=0}^\infty \frac{1}{n!}\] That's why I said it's less than e earlier.

OpenStudy (anonymous):

yes it is , get fraction sum of subsequences

OpenStudy (perl):

wait, its not in that form. how did you see its slightly larger than e/2 , are you doing an inequality

OpenStudy (kainui):

I just calculated e/2 and compared it to your calculation. I knew it had to be less than e because the denominator is the same but contains more terms.

OpenStudy (anonymous):

1/sum(k!) < 1/2 (n)! for all n>2

OpenStudy (anonymous):

so clear sum (sum (k!) < 1/2 sum (1/n!) for n>2 sum(sum k!) <e/2

OpenStudy (perl):

i dont know if this helps http://www.wolframalpha.com/input/?i=1%2Fsum%28k!%2Ck%3D1%2Cn%29+

OpenStudy (perl):

@Marki do you have a typo sum (sum k! ) < e/2 , do you mean sum ( 1/ sum k! ) < e2, and i think that is false e/2 < sum (1 / sum k! ) < e

OpenStudy (perl):

@Marki do you have a typo you wrote sum (sum k! ) < e/2 , did you mean sum ( 1/ sum k! ) < e/2, and i think that is false e/2 < sum (1 / sum k! ) < e

OpenStudy (kainui):

I don't think the fact that it is greater than e/2 is at all significant.

OpenStudy (perl):

right, you can probably find a tighter bound

OpenStudy (kainui):

You might be able to play around this when n gets infinite they approach the same number. \[(1+\frac{1}{n})(1+\frac{1}{n})...(1+\frac{1}{n})=(1+\frac{1}{n})^n=\frac{1}{0!}+\frac{1}{1!}+...+\frac{1}{n!}\]

OpenStudy (perl):

my original impulse was to try and turn the sum into a product , too

OpenStudy (anonymous):

its all about comparison , as u can see that our sequence is less than 1/2 sum (1/n!) the second one is converge thus the first one converge , however im trying to compute how to do it for infinite can't see anything with fourier xD if you have a magical function please show it :)

OpenStudy (kainui):

Another fun fact, the number is between cosh(1) and sinh(1). I say this because \[\large \cosh(1)= \frac{1}{0!}+ \frac{1}{2!}+ \frac{1}{4!}+...\]

OpenStudy (anonymous):

and here is what i was trying to show u guys